# AutoScale: Automatic Prediction of Compute-optimal Data Composition for Training LLMs

### v2

@@ C.1 Theorem 1: Scaling Law for Optimal Data Compositions
$\min_{\mathbf{N}}\left\{\sum_{i=1}^{m}\beta_{i}N_{i}^{-\gamma_{i}}\Bigg{|}\sum_{i=1}^{m}N_{i}=N\right\}$
For any two compute budgets $N^{(1)}\neq N^{(2)}$, let $\mathbf{N}^{(1)*}$ and $\mathbf{N}^{(2)*}$ be their respective minimizers. Then, for any third compute budget $N^{(3)}$ such that $N^{(1)}\neq N^{(3)}\neq N^{(2)}$, the corresponding minimizer $\mathbf{N}^{(3)*}$ must satisfy $\mathbf{N}^{(3)*}=\mathbf{N}^{(2)*}(\mathbf{N}^{(1)*})^{-1}\mathbf{N}^{(2)*}$.
@@
For the evaluation loss given as
$\mathcal{L}=\sum_{i=1}^{m}\beta_{i}\cdot N_{i}^{-\gamma_{i}},$
at an optimal data composition, KKT condition [Bertsekas, 1997] gives that we have the partial derivative of the loss function w.r.t. the amount of data from each domain equal to each other. This gives, for any two domains $a$ and $b$ (w.l.o.g, we simplify the derivation to the case of two domains) with optimal data quantity $N_{a}^{*}$ and $N_{b}^{*}$, respectively, we have
$\frac{\partial\mathcal{L}}{\partial N_{a}}=-\beta_{a}\cdot\gamma_{a}\cdot N_{a}^{-\gamma_{a}-1}$
$\frac{\partial\mathcal{L}}{\partial N_{b}}=-\beta_{b}\cdot\gamma_{b}\cdot N_{b}^{-\gamma_{b}-1}$
$\left.\frac{\partial\mathcal{L}}{\partial N_{a}}\right|_{N_{a}=N_{a}^{*}}=\left.\frac{\partial\mathcal{L}}{\partial N_{b}}\right|_{N_{b}=N_{b}^{*}}$
With straightforward derivations, this gives
$-\beta_{a}\cdot\gamma_{a}\cdot(N_{a}^{*})^{-\gamma_{a}-1}$ $=-\beta_{b}\cdot\gamma_{b}\cdot(N_{b}^{*})^{-\gamma_{b}-1}$ (4)
$\frac{\beta_{a}\cdot\gamma_{a}}{\beta_{b}\cdot\gamma_{b}}$ $=\frac{(N_{a}^{*})^{\gamma_{a}+1}}{(N_{b}^{*})^{\gamma_{b}+1}}$
$(N_{a}^{*})^{\gamma_{a}+1}$ $=\frac{\beta_{a}\gamma_{a}}{\beta_{b}\gamma_{b}}(N_{b}^{*})^{\gamma_{b}+1}$
$N_{a}^{*}$ $=\left[\frac{\beta_{a}\gamma_{a}}{\beta_{b}\gamma_{b}}(N_{b}^{*})^{\gamma_{b}+1}\right]^{\frac{1}{\gamma_{a}+1}}$
Let $N_{a}^{(2)*},N_{b}^{(2)*}$ be the optimal data quantity for domains $a$ and $b$ at a different data scale $N^{(2)}=N_{a}^{(2)*}+N_{b}^{(2)*}\neq N$. Assuming we have the optimal data quantity for domain $b$ becoming $m$ times than $N_{b}^{*}$, namely,
$N_{b}^{(2)*}$ $:=m\cdot N_{b}^{*}$
Then, from Eq. 4, the optimal data quantity for domain $a$ can be given as
$N_{a}^{(2)*}$ $=\left[\frac{\beta_{a}\gamma_{a}}{\beta_{b}\gamma_{b}}(N_{b}^{(2)*})^{\gamma_{b}+1}\right]^{\frac{1}{\gamma_{a}+1}}$ (5)
$=\left[\frac{\beta_{a}\gamma_{a}}{\beta_{b}\gamma_{b}}(m\cdot N_{b}^{*})^{\gamma_{b}+1}\right]^{\frac{1}{\gamma_{a}+1}}$
$=m^{\frac{\gamma_{b}+1}{\gamma_{a}+1}}\cdot\left[\frac{\beta_{a}\gamma_{a}}{\beta_{b}\gamma_{b}}(N_{b}^{*})^{\gamma_{b}+1}\right]^{\frac{1}{\gamma_{a}+1}}$
$=m^{\frac{\gamma_{b}+1}{\gamma_{a}+1}}\cdot N_{a}^{*}$
It can be immediately seen that the optimal domain data $N_{a}^{*}$ and $N_{b}^{*}$ scale at different rates–new optimal data quantity for domain $a$ does not become $m$ times than before. This implies that the optimal data composition is scale-dependent and is different for different training data scales. This implies that the optimal data composition is scale-dependent and is different for different training data scales, establishing the main argument of this paper.
Since the ratio from Eq. 5, $(\gamma_{b}+1)/(\gamma_{a}+1)$, is constant and invariant to the change in the training data scale, it can be utilized to provide a direct approach for predicting the scaling of optimal data compositions–given as
$N_{a}^{(2)*}=(\frac{N_{b}^{(2)*}}{N_{b}^{*}})^{{\frac{\gamma_{b}+1}{\gamma_{a}+1}}}N_{a}^{*}$
Equivalently, taking the logarithm for both sides of the equation, we have
$\log N_{a}^{(2)*}=\log(\frac{N_{b}^{(2)*}}{N_{b}^{*}})^{{\frac{\gamma_{b}+1}{\gamma_{a}+1}}}+\log N_{a}^{*}$
Further, we show that one does not need to estimate any of the coefficients ($\gamma_{a}$, $\gamma_{b}$) from the loss function to predict the optimal data quantity for each domain. Assume one have obtained the optimal data quantity for domains $a$ and $b$, $N_{a}^{(1)*},N_{b}^{(1)*}$, at a data scale $N^{(1)}=N_{a}^{(1)*}+N_{b}^{(1)*}$ and the optimal data quantity $N_{a}^{(2)*},N_{b}^{(2)*}$ at a data scale $N^{(2)}=N_{a}^{(2)*}+N_{b}^{(2)*}$ where $N^{(1)}\neq N^{(2)}$. This gives
$\log N_{a}^{(2)*}=\frac{\gamma_{b}+1}{\gamma_{a}+1}\cdot(\log N_{b}^{(2)*}-\log N_{b}^{(1)*})+\log N_{a}^{(1)*}$
Then, for a different data scale where the optimal data quantity for domain $b$ is $N_{b}^{(3)*}$, the optimal data quantity for domain $a$ can be given as
$\log N_{a}^{(3)*}=$ $\frac{\gamma_{b}+1}{\gamma_{a}+1}\cdot(\log N_{b}^{(3)*}-\log N_{b}^{(2)*})+\log N_{a}^{(2)*}$
$=$ $\frac{\log N_{a}^{(2)*}-\log N_{a}^{(1)*}}{\log N_{b}^{(2)*}-\log N_{b}^{(1)*}}\cdot(\log N_{b}^{(3)*}-\log N_{b}^{(2)*})+\log N_{a}^{(2)*}$
W.l.o.g., consider $\frac{N_{b}^{(3)*}}{N_{b}^{(2)*}}=\frac{N_{b}^{(2)*}}{N_{b}^{(1)*}}$ where $\log N_{b}^{(2)*}-\log N_{b}^{(1)*}=\log N_{b}^{(3)*}-\log N_{b}^{(2)*}$, the equation above can be simplified to
$\log N_{a}^{(3)*}=2\log N_{a}^{(2)*}-\log N_{a}^{(1)*}$
and equivalently,
$N_{a}^{(3)*}=\frac{(N_{a}^{(2)*})^{2}}{N_{a}^{(1)*}}$
Defining compact representations $\mathbf{N^{(i)*}}=diag\{N_{a}^{(i)*},N_{b}^{(i)*}\}$, the above results can be written as
$\mathbf{N^{(3)*}}=\mathbf{N^{(2)*}}(\mathbf{N^{(1)*}})^{-1}\mathbf{N^{(2)*}}$
@@
The process can be iterated (e.g., replacing $\mathbf{N^{(1)*}}$ or $\mathbf{N^{(2)*}}$ with $\mathbf{N^{(3)*}}$) to obtain optimal domain data quantity for different data scales. The example below provides a straightforward look on how this result can be operationalized. ∎
Remark 1 (An example).
@@
C.2 Scaling Latent Skills
We extend this theory to a general case where the evaluation loss is the perplexity averaged over training domains. Consider the evaluation is composed of a number of indepedent sub-tasks ("latent skills" [Tiong et al., ]) which are hidden variables, where each of them observes a power law scaling law relationship with the amount of data contributing to this task ("equivalent data size"), $\mathcal{L}=\ell_{0}+\beta_{a}\cdot K_{a}^{-\gamma_{a}}+\beta_{b}\cdot K_{b}^{-\gamma_{b}}+\beta_{c}\cdot K_{c}^{-\gamma_{c}}+\cdots$ where scalar $K_{j}\geq 0$ denote equivalent data size for skillj, and constants $(\beta_{j},\gamma_{j})\geq 0$ are coefficients associated with skillj, respectively. Mathematically, these latent skills can be seen as an orthogonal basis that spans the space of evaluation loss.
Consider training data from each domain $D_{i}$ contributes to these skills to varying degrees, where Equivalent data size for skillj, $K_{j}$, is given as $K_{j}=c_{j,1}\cdot N_{1}+c_{j,2}\cdot N_{2}+\cdots$ where $N_{i}=w_{i}\cdot N$ denotes the amount of training data from domain $D_{i}$ and constant $c_{j,i}$ is the coefficient measuring the degree of contribution between domain $D_{i}$ and skillj. Defining diagonal matrices for training data composition $\mathbf{N}=diag\{N_{1},N_{2},\cdots\}$ and skill data composition $\mathbf{K}=diag\{K_{a},K_{b},\cdots\}$, we have $\mathbf{K}=\mathbf{A}\mathbf{N}$, where $\mathbf{A}_{ji}=c_{j,i}$ is the matrix for coefficients. For simplicity, we consider training data from each domain will be distributed to the skills such that $\forall i,\sum_{j}N_{i}=1$. This gives the amount of total training data from all domains is identical to the amount of total equivalent data for all skills, $\sum_{j}K_{j}=\sum_{i}N_{i}$. For a training data scale $N=\sum_{i}N_{i}=\sum_{j}K_{j}$, define optimal skill data composition $\mathbf{K^{*}}=diag\{K_{a}^{*},K_{b}^{*},\cdots\}$ as the minimizer of $\mathcal{L}$, given as $\mathbf{K^{*}}={\arg\min}_{\sum_{j}K_{j}=N}\ell_{0}+\beta_{a}\cdot K_{a}^{-\gamma_{a}}+\beta_{b}\cdot K_{b}^{-\gamma_{b}}+\cdots$. Theoretically, there can be an infinite number of latent skills. For analysis, we consider a finite number of $k$ independent skills most important for the evaluation. This can considered as performing Principal Components Analysis (PCA) with orthogonal transformation and selecting the first $k$ independent components. We consider the standard scenario with an equal number of relevant skills and data domains where $k=m$ and $\mathbf{A}$ is a square matrix with full rank. This describes the case where this optimization problem is well-defined. We discuss in App. C.2 what will happen in other scenarios. In this case, $\mathbf{A}$ is invertible and the corresponding optimal training data composition for $\mathbf{K^{*}}$ can be given as $\mathbf{N^{*}}=\mathbf{A}^{-1}\mathbf{K^{*}}$.
@@ C.2 Scaling Latent Skills
Consider the evaluation is composed of a number of indepedent sub-tasks ("latent skills") where each of them observes a power law scaling law relationship with the amount of data contributing to this task ("equivalent data size"). Namely, $\mathcal{L}=\ell_{0}+\beta_{a}\cdot K_{a}^{-\gamma_{a}}+\beta_{b}\cdot K_{b}^{-\gamma_{b}}+\beta_{c}\cdot K_{c}^{-\gamma_{c}}+\cdots$ where scalar $K_{j}\geq 0$ denote equivalent data size for skillj, and constants $(\beta_{j},\gamma_{j})\geq 0$ are coefficients associated with skillj, respectively. Define diagonal matrices for training data composition $\mathbf{N}=diag\{N_{1},N_{2},\cdots\}$ and skill data composition $\mathbf{K}=diag\{K_{a},K_{b},\cdots\}$. Consider training data from each domain $D_{i}$ contributes to these skills to varying degrees, given as $\mathbf{K}=\mathbf{A}\mathbf{N}$ where $\mathbf{A}_{ji}=c_{j,i}$ is the matrix for coefficients. Assume the amount of total training data from all domains is identical to the amount of total equivalent data for all skills, $\sum_{j}K_{j}=\sum_{i}N_{i}$. Assume there is a finite number of latent skills and data domains and $\mathbf{A}$ is a square matrix with full rank. For a training data scale $N=\sum_{i}N_{i}=\sum_{j}K_{j}$ , define optimal skill data composition $\mathbf{K^{*}}=diag\{K_{a}^{*},K_{b}^{*},\cdots\}$ as the minimizer of $\mathcal{L}$ s.t. $\sum_{j}K_{j}=N$ with corresponding optimal training data composition If we have optimal data compositions $\mathbf{N^{(1)*}}=diag\{N_{a}^{(1)*},N_{b}^{(1)*},\cdots\}$ where its corresponding skill data composition $\mathbf{K^{(1)*}}=diag\{K_{a}^{(1)*},K_{b}^{(1)*},\cdots\}=\mathbf{A}\mathbf{N^{(1)*}}$ minimizes $\mathcal{L}$ s.t. $\sum_{j}K_{j}=\sum_{i}N^{(1)*}=N^{(1)}$, and $\mathbf{N^{(2)*}}=diag\{N_{a}^{(2)*},N_{b}^{(2)*},...\}$ where its corresponding skill data composition $\mathbf{K^{(2)*}}=diag\{K_{a}^{(2)*},K_{b}^{(2)*},...\}=\mathbf{A}\mathbf{N^{(2)*}}$ minimizes $\mathcal{L}$ s.t. $\sum_{j}K_{j}^{(2)*}=\sum_{i}N^{(2)*}=N^{(2)}$ where $N^{(2)}\neq N^{(1)}$, then, other optimal data compositions $\mathbf{N^{(3)*}}=diag\{N_{a}^{(3)*},N_{b}^{(3)*},...\}$ where its corresponding skill data composition $\mathbf{K^{(3)*}}=diag\{K_{a}^{(3)*},K_{b}^{(3)*},\cdots\}=\mathbf{A}\mathbf{N^{(3)*}}$ minimizes $\mathcal{L}$ s.t. $\sum_{j}K_{j}^{(3)*}=\sum_{i}N^{(3)*}=N^{(3)}$ where $N^{(3)}\neq N^{(2)}\neq N^{(1)}$ can be given as $\mathbf{N^{(3)*}}=\mathbf{N^{(2)*}}(\mathbf{N^{(1)*}})^{-1}\mathbf{N^{(2)*}}$.
@@
$\mathbf{A}\mathbf{N^{(1)*}}=\mathbf{K^{(1)*}},\quad\mathbf{A}\mathbf{N^{(2)*}}=\mathbf{K^{(2)*}},\quad\mathbf{A}\mathbf{N^{(3)*}}=\mathbf{K^{(3)*}}$
From results of Theorem 1, we have
$\mathbf{K^{(3)*}}=\mathbf{K^{(2)*}}(\mathbf{K}^{(1)*})^{-1}\mathbf{K^{(2)*}}$
which gives
$\mathbf{A}\mathbf{N^{(3)*}}=(\mathbf{A}\mathbf{N^{(2)*}})(\mathbf{A}\mathbf{N^{(1)*}})^{-1}\mathbf{A}\mathbf{N^{(2)*}}$
@@
$(\mathbf{A}\mathbf{N^{(1)*}})^{-1}=(\mathbf{N^{(1)*}})^{-1}\mathbf{A}^{-1}$
@@
$\mathbf{A}\mathbf{N^{(3)*}}=\mathbf{A}\mathbf{N^{(2)*}}[(\mathbf{N^{(1)*}})^{-1}\mathbf{A}^{-1}]\mathbf{A}\mathbf{N^{(2)*}}=\mathbf{A}\mathbf{N^{(2)*}}(\mathbf{N^{(1)*}})^{-1}\mathbf{N^{(2)*}}$

### v5

@@ B.1 Theorem 1: Scale-Dependent Optimal Composition
$\min_{\mathbf{N}}\left\{\sum_{i=1}^{m}\beta_{i}N_{i}^{-\gamma_{i}}\Bigg|\sum_{i=1}^{m}N_{i}=N\right\}.$
For any two compute budgets $N^{(1)}\neq N^{(2)}$, let $\mathbf{N}^{*}(N^{(1)})$ and $\mathbf{N}^{*}(N^{(2)})$ be their respective minimizers. For any third data composition $\mathbf{N}(N^{(3)})$, if there exists some constant $k\in\mathbb{R}^{+}$ such that
$\mathbf{N}(N^{(3)})=\mathbf{N}^{*}(N^{(2)})[(\mathbf{N}^{*}(N^{(1)}))^{-1}\mathbf{N}^{*}(N^{(2)})]^{k},$
then, $\mathbf{N}(N^{(3)})$ is the minimizer for data budget $N^{(3)}=\sum_{i=1}^{m}N^{(3)}_{i}$, given as
$\mathbf{N}(N^{(3)})=\arg\min_{\mathbf{N}}\left\{\sum_{i=1}^{m}\beta_{i}N_{i}^{-\gamma_{i}}\Bigg|\sum_{i=1}^{m}N_{i}=N^{(3)}\right\}=\mathbf{N}^{*}(N^{(3)}).$
@@
Setup: We begin with the following optimization problem, defined at a given total training data scale $N^{(1)}$:
$\min_{\mathbf{N}}\left\{\sum_{i=1}^{m}\beta_{i}N_{i}^{-\gamma_{i}}\,\middle|\,\sum_{i=1}^{m}N_{i}=N^{(1)}\right\}.$
Here, $\mathbf{N}=\mathrm{diag}\{N_{1},N_{2},\ldots,N_{m}\}$ is a diagonal matrix whose diagonal entries are the amounts of data allocated to each of the $m$ domains.
For this problem, there exists a unique optimal solution $\mathbf{N}^{*}(N^{(1)})=\mathrm{diag}\{N_{1}^{(1)*},N_{2}^{(1)*},\ldots,N_{m}^{(1)*}\}$. This $\mathbf{N}^{*}(N^{(1)})$ represents the compute-optimal data composition at the data scale $N^{(1)}$.
First-Order Conditions (KKT): At optimality, the Karush–Kuhn–Tucker (KKT) conditions ensure that the partial derivatives of the objective function with respect to each $N_{i}$ are equal (up to the same Lagrange multiplier for the equality constraint $\sum_{i}N_{i}=N^{(1)}$). For any pair of domains $a$ and $b$, we must have:
$\left.\frac{\partial}{\partial N_{a}}\left(\sum_{i=1}^{m}\beta_{i}N_{i}^{-\gamma_{i}}\right)\right|_{N_{a}=N_{a}^{(1)*}}=\left.\frac{\partial}{\partial N_{b}}\left(\sum_{i=1}^{m}\beta_{i}N_{i}^{-\gamma_{i}}\right)\right|_{N_{b}=N_{b}^{(1)*}}.$
Computing these derivatives, we get:
$\beta_{a}\gamma_{a}(N_{a}^{(1)*})^{-\gamma_{a}-1}=-\beta_{b}\gamma_{b}(N_{b}^{(1)*})^{-\gamma_{b}-1}.$
From this equality:
$\frac{\beta_{a}\gamma_{a}}{\beta_{b}\gamma_{b}}=\frac{(N_{a}^{(1)*})^{\gamma_{a}+1}}{(N_{b}^{(1)*})^{\gamma_{b}+1}}.$
Rearranging, we obtain a fundamental scaling relationship:
$N_{a}^{(1)*}=\left(\frac{\beta_{a}\gamma_{a}}{\beta_{b}\gamma_{b}}(N_{b}^{(1)*})^{\gamma_{b}+1}\right)^{\frac{1}{\gamma_{a}+1}}.$
Scaling to a Second Data Scale $N^{(2)}$: Now consider a different total data scale $N^{(2)}\neq N^{(1)}$, with the corresponding compute-optimal solution $\mathbf{N}^{*}(N^{(2)})=\mathrm{diag}\{N_{1}^{(2)*},N_{2}^{(2)*},\ldots,N_{m}^{(2)*}\}$.
Suppose we know how the optimal amount of data for domain $b$ changes from $N^{(1)}$ to $N^{(2)}$. Specifically, let:
$N_{b}^{(2)*}=m\cdot N_{b}^{(1)*}$
for some scaling factor $m>0$.
Applying the same relationship used for the first scale, but now at the second scale, we find that for domain $a$:
$N_{a}^{(2)*}=\left(\frac{\beta_{a}\gamma_{a}}{\beta_{b}\gamma_{b}}(N_{b}^{(2)*})^{\gamma_{b}+1}\right)^{\frac{1}{\gamma_{a}+1}}=\left(\frac{\beta_{a}\gamma_{a}}{\beta_{b}\gamma_{b}}(m\cdot N_{b}^{(1)*})^{\gamma_{b}+1}\right)^{\frac{1}{\gamma_{a}+1}}.$
This simplifies to:
$N_{a}^{(2)*}=m^{\frac{\gamma_{b}+1}{\gamma_{a}+1}}N_{a}^{(1)*}.$ (2)
Notice that $m^{\frac{\gamma_{b}+1}{\gamma_{a}+1}}\neq m$ in general. Thus, when the budget scales by a factor $m$ in domain $b$, the optimal amount for domain $a$ scales by a different factor. This shows that the optimal composition is scale-dependent.
Predicting a Third Scale $N^{(3)}$: We now know the optimal compositions at two scales $N^{(1)}$ and $N^{(2)}$. Consider a third scale $N^{(3)}$ and its optimal solution $\mathbf{N}^{*}(N^{(3)})=\mathrm{diag}\{N_{1}^{(3)*},N_{2}^{(3)*},\ldots,N_{m}^{(3)*}\}$.
If we choose $N_{b}^{(3)*}$ such that:
$\frac{N_{b}^{(3)*}}{N_{b}^{(2)*}}=\frac{N_{b}^{(2)*}}{N_{b}^{(1)*}},$ (3)
then the change in $N_{b}$ from $N^{(2)}$ to $N^{(3)}$ mirrors the change from $N^{(1)}$ to $N^{(2)}$.
Since the scaling exponent $\frac{\gamma_{b}+1}{\gamma_{a}+1}$ remains the same, this symmetrical setup leads to:
$N_{a}^{(3)*}=\frac{(N_{a}^{(2)*})^{2}}{N_{a}^{(1)*}}.$
Matrix Form: Because all domains scale in a similar fashion, we can write this relationship compactly using diagonal matrices. Define:
$\mathbf{N^{*}}(N^{(i)})=\mathrm{diag}\{N_{1}^{(i)*},N_{2}^{(i)*},\ldots,N_{m}^{(i)*}\}.$
The element-wise relationship $\frac{(N_{a}^{(2)*})^{2}}{N_{a}^{(1)*}}$ for each domain $a$ can be expressed as:
$\mathbf{N}^{*}(N^{(3)})=\mathbf{N}^{*}(N^{(2)})(\mathbf{N}^{*}(N^{(1)}))^{-1}\mathbf{N}^{*}(N^{(2)}).$
Here, $(\mathbf{N}^{*}(N^{(1)}))^{-1}$ is the inverse of the diagonal matrix $\mathbf{N}^{*}(N^{(1)})$, obtained by taking the reciprocal of each diagonal element $N_{a}^{(1)*}$.
We have shown that given two distinct data scales $N^{(1)}$ and $N^{(2)}$ and their corresponding optimal solutions $\mathbf{N}^{*}(N^{(1)})$ and $\mathbf{N}^{*}(N^{(2)})$, one can construct a third optimal solution $\mathbf{N}^{*}(N^{(3)})$ using the formula:
$\mathbf{N}^{*}(N^{(3)})=\mathbf{N}^{*}(N^{(2)})(\mathbf{N}^{*}(N^{(1)}))^{-1}\mathbf{N}^{*}(N^{(2)}).$
This relationship holds without needing to explicitly estimate the parameters $\gamma_{i}$ or $\beta_{i}$, and it confirms that the optimal data composition is scale-dependent. Thus, the given scaling law for optimal data compositions is established.
Generalization to Prediction for Any Data Scale: Finally, we generalize from the case in Eq. (3) to allow prediction of optimal data composition for any data scale. Consider for some constant $\forall k\in\mathbb{R}^{+}$, we choose $N_{b}^{(k)*}$ such that
$\frac{N_{b}^{(k)*}}{N_{b}^{(2)*}}=\left(\frac{N_{b}^{(2)*}}{N_{b}^{(1)*}}\right)^{k}=m^{k},$
Same as the procedure in Eq. (2), KKT optimality conditions yield the corresponding optimal data quantity for domain a at the same scale as $N_{b}^{(k)*}$ as
$N_{a}^{(k)*}=(m^{k})^{\frac{\gamma_{b}+1}{\gamma_{a}+1}}N_{a}^{(2)*}=\left(\frac{N_{a}^{(2)*}}{N_{a}^{(1)*}}\right)^{k}N_{a}^{(2)*}.$
Rearranging in the matrix form, we have the following formula
$\mathbf{N}^{*}(N^{(k)})=\mathbf{N}^{*}(N^{(2)})[\mathbf{N}^{*}(N^{(2)})(\mathbf{N}^{*}(N^{(1)}))^{-1}]^{k},$
@@
Application in AutoScale: Note that this formulate holds for any $k\in\mathbb{R}^{+}$. Thus, by scanning through the values of $k$, one can find optimal data composition $\mathbf{N}^{*}$ for any target data scale $N=\sum_{i}N^{*}_{i}$. In practice, for a target data scale $N>N^{(2)}=\sum_{i}N^{(2)*}_{i}$, one only needs to conduct a line search along $k>1$ to find the value of $k$ where $\sum_{i}N^{(k)*}_{i}=N$ to determine its corresponding optimal data composition $\mathbf{N}^{*}$.
Remark B.2 (An example).
@@
B.2 Scaling Latent Skills
We extend this theory to a general case where the evaluation loss is the perplexity averaged over training domains. Consider the evaluation is composed of a number of independent sub-tasks ("latent skills" (Tiong et al., )) which are hidden variables, where each of them observes a power law scaling law relationship with the amount of data contributing to this task ("equivalent data size"), $\mathcal{L}=\ell_{0}+\beta_{a}\cdot K_{a}^{-\gamma_{a}}+\beta_{b}\cdot K_{b}^{-\gamma_{b}}+\beta_{c}\cdot K_{c}^{-\gamma_{c}}+\cdots$ where scalar $K_{j}\geq 0$ denote equivalent data size for skillj, and constants $(\beta_{j},\gamma_{j})\geq 0$ are coefficients associated with skillj, respectively. Mathematically, these latent skills can be seen as an orthogonal basis that spans the space of evaluation loss.
Consider training data from each domain $D_{i}$ contributes to these skills to varying degrees, where Equivalent data size for skillj, $K_{j}$, is given as $K_{j}=c_{j,1}\cdot N_{1}+c_{j,2}\cdot N_{2}+\cdots$ where $N_{i}=w_{i}\cdot N$ denotes the amount of training data from domain $D_{i}$ and constant $c_{j,i}$ is the coefficient measuring the degree of contribution between domain $D_{i}$ and skillj. Defining diagonal matrices for training data composition $\mathbf{N}=diag\{N_{1},N_{2},\cdots\}$ and skill data composition $\mathbf{K}=diag\{K_{a},K_{b},\cdots\}$, we have $\mathbf{K}=\mathbf{A}\mathbf{N}$, where $\mathbf{A}_{ji}=c_{j,i}$ is the matrix for coefficients. For simplicity, we consider training data from each domain will be distributed to the skills such that $\forall i,\sum_{j}N_{i}=1$. This gives the amount of total training data from all domains is identical to the amount of total equivalent data for all skills, $\sum_{j}K_{j}=\sum_{i}N_{i}$. For a training data scale $N=\sum_{i}N_{i}=\sum_{j}K_{j}$, define optimal skill data composition $\mathbf{K^{*}}=diag\{K_{a}^{*},K_{b}^{*},\cdots\}$ as the minimizer of $\mathcal{L}$, given as $\mathbf{K^{*}}={\arg\min}_{\sum_{j}K_{j}=N}\ell_{0}+\beta_{a}\cdot K_{a}^{-\gamma_{a}}+\beta_{b}\cdot K_{b}^{-\gamma_{b}}+\cdots$. Theoretically, there can be an infinite number of latent skills. For analysis, we consider a finite number of $k$ independent skills most important for the evaluation. This can considered as performing Principal Components Analysis (PCA) with orthogonal transformation and selecting the first $k$ independent components. We consider the standard scenario with an equal number of relevant skills and data domains where $k=m$ and $\mathbf{A}$ is a square matrix with full rank. This describes the case where this optimization problem is well-defined. We discuss in App. B.2 what will happen in other scenarios. In this case, $\mathbf{A}$ is invertible and the corresponding optimal training data composition for $\mathbf{K^{*}}$ can be given as $\mathbf{N^{*}}=\mathbf{A}^{-1}\mathbf{K^{*}}$.
@@ B.2 Scaling Latent Skills
Consider the evaluation is composed of a number of independent sub-tasks ("latent skills") where each of them observes a power law scaling law relationship with the amount of data contributing to this task ("equivalent data size"). Namely,
$\mathcal{L}=\ell_{0}+\beta_{a}\cdot K_{a}^{-\gamma_{a}}+\beta_{b}\cdot K_{b}^{-\gamma_{b}}+\beta_{c}\cdot K_{c}^{-\gamma_{c}}+\cdots$
where scalar $K_{j}\geq 0$ denote equivalent data size for skillj, and constants $(\beta_{j},\gamma_{j})\geq 0$ are coefficients associated with skillj, respectively. Define diagonal matrices for training data composition $\mathbf{N}=diag\{N_{1},N_{2},\cdots\}$ and skill data composition $\mathbf{K}=diag\{K_{a},K_{b},\cdots\}$. Consider training data from each domain $D_{i}$ contributes to these skills to varying degrees, given as $\mathbf{K}=\mathbf{A}\mathbf{N}$ where $\mathbf{A}_{ji}=c_{j,i}$ is the matrix for coefficients. Assume the amount of total training data from all domains is identical to the amount of total equivalent data for all skills, $\sum_{j}K_{j}=\sum_{i}N_{i}$. Assume there is a finite number of latent skills and data domains and $\mathbf{A}$ is a square matrix with full rank.
For a training data scale $N=\sum_{i}N_{i}=\sum_{j}K_{j}$ , define optimal skill data composition $\mathbf{K^{*}}=diag\{K_{a}^{*},K_{b}^{*},\cdots\}$ as the minimizer of $\mathcal{L}$ s.t. $\sum_{j}K_{j}=N$ with corresponding optimal training data composition. If we have optimal data compositions $\mathbf{N^{*}}(N^{(1)})=diag\{N_{a}^{(1)*},N_{b}^{(1)*},\cdots\}$ where its corresponding skill data composition $\mathbf{K^{(1)*}}=diag\{K_{a}^{(1)*},K_{b}^{(1)*},\cdots\}=\mathbf{A}\mathbf{N^{*}}(N^{(1)})$ minimizes $\mathcal{L}$ s.t. $\sum_{j}K_{j}=\sum_{i}N^{(1)*}=N^{(1)}$, and $\mathbf{N^{*}}(N^{(2)})=diag\{N_{a}^{(2)*},N_{b}^{(2)*},$...\} where its corresponding skill data composition $\mathbf{K^{(2)*}}=diag\{K_{a}^{(2)*},K_{b}^{(2)*},...\}=\mathbf{A}\mathbf{N^{*}}(N^{(2)})$ minimizes $\mathcal{L}$ s.t. $\sum_{j}K_{j}^{(2)*}=\sum_{i}N^{(2)*}=N^{(2)}$ where $N^{(2)}\neq N^{(1)}$, then, other optimal data compositions $\mathbf{N^{*}}(N^{(3)})=diag\{N_{a}^{(3)*},N_{b}^{(3)*},$...\} where the corresponding skill data composition $\mathbf{K^{(3)*}}=diag\{K_{a}^{(3)*},K_{b}^{(3)*},\cdots\}=\mathbf{A}\mathbf{N^{*}}(N^{(3)})$ minimizes $\mathcal{L}$ s.t. $\sum_{j}K_{j}^{(3)*}=\sum_{i}N^{(3)*}=N^{(3)}$ where $N^{(3)}\neq N^{(2)}\neq N^{(1)}$ must satisfy
$\mathbf{N^{*}}(N^{(3)})=\mathbf{N^{*}}(N^{(2)})[(\mathbf{N^{*}}(N^{(1)}))^{-1}\mathbf{N^{*}}(N^{(2)})]^{k}$
for some $k\in\mathbb{R}^{+}$.
@@
$\mathbf{A}\mathbf{N^{*}}(N^{(1)})=\mathbf{K^{(1)*}},\quad\mathbf{A}\mathbf{N^{*}}(N^{(2)})=\mathbf{K^{(2)*}},\quad\mathbf{A}\mathbf{N^{*}}(N^{(3)})=\mathbf{K^{(3)*}}$
From results of Theorem 1 in Section 3.2.2, we have
$\mathbf{K^{(3)*}}=\mathbf{K^{(2)*}}[(\mathbf{K}^{(1)*})^{-1}\mathbf{K^{(2)*}}]^{k}$
for some $k\in\mathbb{R}^{+}$, which gives
$\mathbf{A}\mathbf{N^{*}}(N^{(3)})=(\mathbf{A}\mathbf{N^{*}}(N^{(2)}))[(\mathbf{A}\mathbf{N^{*}}(N^{(1)}))^{-1}\mathbf{A}\mathbf{N^{*}}(N^{(2)})]^{k}$
@@
$(\mathbf{A}\mathbf{N^{*}}(N^{(1)}))^{-1}=(\mathbf{N^{*}}(N^{(1)}))^{-1}\mathbf{A}^{-1}$
@@
$\mathbf{A}\mathbf{N^{*}}(N^{(3)})=\mathbf{A}\mathbf{N^{*}}(N^{(2)})[(\mathbf{N^{*}}(N^{(1)}))^{-1}\mathbf{A}^{-1}\mathbf{A}\mathbf{N^{*}}(N^{(2)})]^{k}=\mathbf{A}\mathbf{N^{*}}(N^{(2)})[(\mathbf{N^{*}}(N^{(1)}))^{-1}\mathbf{N^{*}}(N^{(2)})]^{k}$